OpenStudy (anonymous):

so someone should teach me a simple way to do half lives, it'd be very helpful!

OpenStudy (goformit100):

What the question suggests ?

OpenStudy (anonymous):

For times that you can divide wholely into...for example Half-life of 60 seconds, find blah blah amount in 3 minutes Just make a table Time Amount Remaining 0 100% 60 sec 50% etc etc For more complicated problems, resort to the exponential decay formula: The half-life formula is Nt = No(1/2)^(T/Tt), where Nt is the quantity that still remains and has not yet decayed No is the initial quantity of the thing that will decay T is the amount of time that has passed Tt is the half-life

OpenStudy (anonymous):

the equations my teacher gave to me were t(1/2)=0.693/K and ln[A]=-Kt+ln[Ao]. How would i use those in a story problem?

OpenStudy (aaronq):

t(1/2)=0.693/K and ln[A]=-Kt+ln[Ao] are equations for a first order reaction calculating the half life from this is easy, just find K notice that ln[A]=-Kt+ln[Ao] looks like y = mx+b